Evaulate the piecewise functions at the indicated values
f(x)=
2x if x < −4
x + 4 if −4 ≤ x ≤ 3
(x − 3)2 if x > 3
f(-5)=
f(-4)
f(2)=
f(3)=
f(6)=

Answer by stanbon(75887)

Evaulate the piecewise functions at the indicated values
f(x)=
2x if x < −4
x + 4 if −4 ≤ x ≤ 3
(x − 3)2 if x > 3
------------------
f(-5)= 2*-5 = -10
-----------------------
f(-4) = -4+4 = 0
--------------------
f(2) = 2+4 = 6
------------------------
f(3) = 3+4 = 7
------------------------------
f(6) = 6-3 = 3
===================
Cheers,
Stan H.
-------------